OpenStudy (anonymous):

HELPPP!!!! WAT IS 1+1 i forgot how to add

OpenStudy (anonymous):

Kids these days

OpenStudy (shadowlegendx):

11

OpenStudy (anonymous):

^

OpenStudy (anonymous):

so wat is answer

OpenStudy (shadowlegendx):

Answer = 11

OpenStudy (anonymous):

me 2 :(

OpenStudy (anonymous):

kids these days or 11

OpenStudy (anonymous):

2

OpenStudy (anonymous):

lol @RachelKate SAME :D

OpenStudy (anonymous):

are you serious.... this question is wasting peoples time and space.

OpenStudy (anonymous):

73465548489 thats the answer i promise ;)

OpenStudy (solomonzelman):

Looks like \(\normalsize\color{blue}{ 1+1i }\) (looking at the original question) you can not simplify that more.

OpenStudy (anonymous):

2 lol

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 1+1i }\) can be re-written as \(\normalsize\color{blue}{ 1+i }\)

OpenStudy (shadowlegendx):

@_angeliquexoxo people can choose to waste their time were, as for space, this question will soon go under the hundreds of math questions that pop up every day

OpenStudy (anonymous):

@_angeliquexoxo JUST YOUR TIME ITS A SERIOUS QUESTION

OpenStudy (shadowlegendx):

here*

OpenStudy (anonymous):

lol here i am to stop the madness i have the answer to all of your problems people

OpenStudy (anonymous):

the answer is 2

OpenStudy (solomonzelman):

Or √2i or 2i^( ½ ) if you want to write it as one term.

OpenStudy (anonymous):

@issaiahphillips THANKS

OpenStudy (solomonzelman):

I should say │ √2i │

OpenStudy (anonymous):

OMG 2 howd u get that???? 0.0

OpenStudy (anonymous):

lol no prob anytime @Kidz_Wavy

OpenStudy (anonymous):

IDK CAN U EXPLAIN HOW U GOT HTAT ANSWER

OpenStudy (anonymous):

PLEASE

OpenStudy (anonymous):

LOL THAT TYPO

OpenStudy (anonymous):

i really need help!!!

OpenStudy (anonymous):

howd i get into 10th grade????

OpenStudy (anonymous):

lol well 1 and then u add a one to that again its 2

OpenStudy (anonymous):

lol like 1 +1+1 is 3

OpenStudy (anonymous):

Im in that same grade and i dont now how to add 1+1

OpenStudy (anonymous):

omg ive been enlightened

OpenStudy (solomonzelman):

I can graph \(\normalsize\color{blue}{ 1+i }\) if you want (imaginary line) xD

OpenStudy (anonymous):

I think i get it 1+1=11

OpenStudy (anonymous):

@SolomonZelman NO but can u teach me how to add

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@Kidz_Wavy i think its 110 not 11 sorry

OpenStudy (anonymous):

@ShadowLegendX excuse me? Your probably just defending this kind of stupidity because you can relate to it. This is a website where people come for help and this moron is sabotaging it. I mean look at @SolomonZelman wasting his time trying to help this kid and this kid is using it as a joke. Immaturity is what it is.

OpenStudy (solomonzelman):

2 years ago in Biology: 1 + 1, when two people (of different genders) so to speak "add together" or are "added together" to form a child, this gives that the addition of these 2 people ( 1 + 1 ) can be equal to 2 parents and 1 child or 2 parents and 2 children (or even three children) This gives that 1 + 1 = 2 or 3 etc., (up to a certain limit of children in one woman of course)

OpenStudy (anonymous):

Ohh i get it @RachelKate 1+1=110

OpenStudy (anonymous):

@Kidz_Wavy yuppp yayy weve figured things out together!!!

OpenStudy (anonymous):

You need the answer and your not being stupid then the answer is 2 for Christ's sake.

OpenStudy (anonymous):

@_angeliquexoxo ur not nice im reporting u

OpenStudy (anonymous):

oh man wow umm i guess i outta go now things look like it is getting pretty steamy up in here and people r getting salty

OpenStudy (solomonzelman):

@_angeliquexoxo, so you do believe in G-d ?

OpenStudy (anonymous):

Yaa @_angeliquexoxo do you do believe in G-d ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

I do @SolomonZelman

OpenStudy (anonymous):

@issaiahphillips k thanks

OpenStudy (thomas5267):

You must believe in the power of GCD lol.

OpenStudy (anonymous):

@_angeliquexoxo can u prove it

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

@thomas5267 LMAO

OpenStudy (anonymous):

haha i am lloving this conversation

OpenStudy (solomonzelman):

No one can prove that God exists or the opposite, it is a question of faith:) Lets get back to our unsolved question , \(\normalsize\color{blue}{ 1+1 }\) or \(\normalsize\color{blue}{ 1+1i }\) what ever it was....

OpenStudy (anonymous):

@_angeliquexoxo state the 10 commandments

OpenStudy (thomas5267):

Are you working on the field with two elements? If so, 1+1=0 !

OpenStudy (anonymous):

So wat is the answer to this question 1+1 is it 11 110 0 five or GOD

OpenStudy (solomonzelman):

@Kidz_Wavy don't know why people call them 10 commandments if there are more than 10 with in them. There are 13 actually. That's what I (as an Orthodox Jew) believe.

OpenStudy (anonymous):

@SolomonZelman CAN U STATE THEM PLZ AND ANWER THIS QUESTION 1+1

OpenStudy (anonymous):

lol this is still going on

OpenStudy (anonymous):

YES AND UR APART OF IT NOW @alenahall12 HELP WAT IS 1+1 :D

OpenStudy (thomas5267):

If you are working on the trivial ring, 1+1=1. If you are working with binary numbers, 1+1=10. What numbers are you working with?

OpenStudy (anonymous):

oh great lol well lets see if you went to like prek you should know it = two unless you mean it in a different way of asking lol

OpenStudy (anonymous):

Is this the krusty Krab?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so why is this a long list when the answer is simple

OpenStudy (anonymous):

it is 11

OpenStudy (anonymous):

@Gin_Ichimaru yes i think it is because i cant find the answer to this question so wat is 1+1

OpenStudy (anonymous):

i cant either

OpenStudy (anonymous):

i give up on this

OpenStudy (anonymous):

i actually went online to search for 1 + 1

OpenStudy (anonymous):

and its not there

OpenStudy (thomas5267):

@SolomonZelman There are exactly 10 commandments in base 13.

OpenStudy (anonymous):

out all of these comment i still dont know wat the answer is

OpenStudy (anonymous):

2

OpenStudy (anonymous):

1 + 1 lets do the math

OpenStudy (anonymous):

since someone got mad at me for giving an answer

OpenStudy (solomonzelman):

The answer is \[\frac{d}{dx}~2x\]

OpenStudy (anonymous):

r u sure and if u r that can u plz teach me the way of the addition

OpenStudy (solomonzelman):

I am sure :)

OpenStudy (anonymous):

ok 1+1

OpenStudy (anonymous):

HOW CAN YOU EXPLAIN THAT

OpenStudy (anonymous):

use fingers

OpenStudy (anonymous):

1+1 = 2

OpenStudy (solomonzelman):

You have 1 suit, and then you buy another one. How many suits do you have now ? (If you ever wear suits)

OpenStudy (anonymous):

now i have a question

OpenStudy (anonymous):

i have did not buy suits and i have no fingers

OpenStudy (anonymous):

typo have

OpenStudy (thomas5267):

Impressive! Do you type by facerolling on the keyboard?

OpenStudy (anonymous):

i use my 8=D

OpenStudy (solomonzelman):

you can also write the answer as \(\LARGE\color{blue}{ \int_{ }^{ } 0 }\) this is not precise though xD

OpenStudy (solomonzelman):

sorry for being stupid ( * _ - )

OpenStudy (anonymous):

can someone plzzzzzz just teach me how to add and wat is the answer to 1+1